# H2EAL: Hybrid-Bonding Architecture with Hybrid Sparse Attention for Efficient Long-Context LLM Inference

### v1

@@ I Introduction
Large language models (LLMs) have revolutionized various domains in field of generative artificial intelligence (AI) [1]. While deployment on the cloud remains prevalent for mass user access, emerging trends show LLMs are deployed on intelligent edge platforms to address growing demands for localized functionality, personalized interactions, and data privacy [2, 3]. Edge-oriented LLM services prioritize small-batch processing while addressing growing demands for long-context applications such as long document analysis [4], multi-turn dialogue [5], long chain-of-thought reasoning [6]. Figure 1 demonstrates the trend in long-context processing and highlights hardware challenges in KV cache management.
Efficient LLM inference relies on the hardware to process two phases, i.e., prefill and decoding. During the prefill phase, LLM processes the input prompt, leveraging parallel processing to process multiple tokens. In contrast, during the decoding phase, LLM follows auto-regressive generation and only outputs one token at each step, which is more dependent on the memory due to the low data reuse. To meet both demands, near-memory processing (NMP)-based heterogeneous architectures have been proposed [7, 8, 9]. NMP combines standard computing units with memory-integrated accelerators to improve both compute-intensive and memory-heavy operations. However, traditional NMP approaches with in-die DRAM integration suffer from limited logic density and storage capacity [10]. These challenges lead to lower throughput of attention computation and less effective key-value (KV) cache management as context lengths increase, ultimately affecting real-time performance in scenarios with long contexts and low batch sizes.
@@
The recent emerging hybrid bonding (HB) technology provides a promising alternative to conventional NMP architectures [11, 12, 13, 14, 15]. It not only delivers substantial bandwidth with lower power consumption than HBM[13, 12], but also boosts NMP’s computational capacity by allowing processing engines to be implemented on the vertically stacked logic die. However, it exhibits distinctive features that differentiate it from other architectures. HB adopts a distributed-memory design, where the banks in memory dies are not directly interconnected, but connected to logic banks of high bandwidth with separate I/O and control units. This means that each bank operates largely independently, with a relatively limited storage capacity per bank.
Meanwhile, recent advancements in sparse attention methods [16, 17, 18] have shown strong potential for edge inference in long-context scenarios. These approaches significantly reduce KV cache overhead with minimal impact on the model accuracy, becoming particularly beneficial for resource-constrained edge environments. However, previous methods are primarily designed for GPU-like systems, which rely on unified memory and enable seamless data sharing across all processing cores. These methods are not suitable for HB architectures due to the distributed memory. Additionally, different attention sparsity across heads has been observed in [19, 20]. Therefore, allocating a unified sparse pattern for all heads is insufficient to fully leverage the sparsity.
@@ II-A HB Architecture for Edge-side LLM Inference
Deploying LLMs on edge devices for smart homes, autonomous driving, and IoT demands small-scale, real-time tasks to meet latency and power constraints [2, 3]. For long-context LLM inference where models handle extended lengths of sequence, edge devices face a unique challenge since the KV cache size becomes significant [21]. Storing and accessing large KV cache require high memory bandwidth demands, which are difficult for traditional memory architectures, especially for edge devices with limited resources.
Emerging HB technology presents a promising solution by vertically stacking memory dies (e.g., DRAM) on top of logic dies and interconnecting them through high-bandwidth copper (Cu-Cu) direct bonding [11, 12, 13, 14] as illustrated in Figure 3(a). This 3D integration enables substantial bandwidth and low power consumption, making HB a promising alternative to HBM and in-die NMP. However, despite its advantages, HB still faces challenges in managing memory and compute resources that are distributed spatially, especially for long-context inference where the KV cache becomes the dominant. Balancing the computational demands of LLMs with the memory bandwidth available in HB architecture is crucial for edge-side performance.
@@ II-B Sparse Attention Algorithm
Long-context LLM inference increases the computation and memory requirements induced by the KV cache, which becomes the prominent bottleneck. To alleviate this, sparse attention techniques have been proposed to reduce the number of attended tokens during each decoding step. These methods can be divided into static sparse attention (pattern-based) and dynamic sparse attention (input-based). Static sparsity such as the first few or local tokens uses pre-defined attention patterns to limit the number of tokens retained in the KV cache, which directly reduces memory usage but may risk losing important context [22, 23]. In contrast, dynamic sparsity prunes less relevant tokens according to the input-based attention scores, offering better flexibility and utilization of the memory [16, 17, 24].
Hardware-based sparse attention has been explored to reduce memory usage and computational overhead in LLMs [25, 26, 16]. TranCIM [25] adopts fixed-pattern sparsity based on the StreamingLLM algorithm [22], restricting attention to local tokens and reducing data movement but yielding inflexible, suboptimal long-context performance. SADIMM [26] employs top-k attention-based dynamic sparsity, improving adaptability but incurring hardware overhead for token ranking and risking accuracy loss[17, 24]. H2O [16] evicts low-importance tokens yet underuses attention sparsity[19, 27], and remains bottlenecked by GPU memory bandwidth during decoding. As shown in Table I, these studies underscore the challenges in balancing sparse attention efficiency with hardware complexity, latency, and power consumption.
@@
Method TranCIM[25] SADIMM[26] H2O[16] H2EAL (This work)
@@ III-A Basic LLM Implementation on HB Architectures
Apart from regular multi-head attention (MHA), group-query attention (GQA) [28] and multi-query attention (MQA) [29] are proposed to alleviate the memory demand of MHA as illustrated in Figure 3(c). Different attention mechanisms are essentially the same as we focus on the KV head. Specifically, when centering around the KV head, once the KV cache is properly mapped and handled, the remaining weight modules can be allocated to the corresponding group of banks by head parallelism. However, implementing LLMs on the HB architecture on long contexts still faces the following challenges as shown in Figure 2.
@@ III-B Challenges
Challenge 1: Enormous KV cache overhead motivates the algorithm-hardware co-design. When handling long contexts, LLMs suffer from the increasing computation and memory requirements (including both memory capacity and bandwidth) associated with the KV cache, which becomes the prominent bottleneck for long-context LLM inference and exceeds the storage capacity of HB [21] as illustrated in Figure 2. This challenge motivates us to leverage the inherent attention sparsity to compress the KV cache size. However, existing sparse attention methods either suffer from high accuracy degradation [22] or introduce extra computation costs [24, 17]. Meanwhile, recent studies observe that sparsity varies across different heads [19, 20], providing us with larger optimization space for more flexible and extreme KV cache compression. These observations motivate us to co-design algorithm and hardware, including 1) an accurate head-wise sparse attention algorithm to achieve high sparsity without compromising accuracy; and 2) specialized hardware modules to better leverage the sparse attention with heterogeneous sparsity across heads.
Challenge 2: Varying sparsity in different heads leads to workload imbalance between banks. As described above, sparsity varies across different heads, shown in Figure 3(b). Instead of using uniform sparsity across heads, we apply different sparse patterns for different heads. However, such sparsity heterogeneity among heads leads to critical synchronization bottlenecks due to data dependencies - all heads must await completion of the most computationally intensive operation before progressing to subsequent layers. The challenge can be more severe due to the distributed memory architecture of HB architecture [11, 12, 13, 14], which fundamentally differs from conventional unified memory systems. Hence, how to effectively balance workload distribution across memory-distributed banks for the hybrid sparsity pattern is a key question.
@@ IV-A Hybrid Sparse Attention with Algorithm-Hardware Co-Design
Transformer-based LLMs exhibit different sparse patterns in different heads [19, 27, 23, 20, 18]. Figure 4 shows the sparse patterns varies across different heads and layers. As shown in challenge 1 in Figure 2, some heads dynamically attend to the tokens based the the query while other heads mainly attend to the local and the first few tokens (termed as sink tokens) as observed in [19]. However, existing works set uniform sparse ratio for different heads, which is not sub-optimal for both efficiency and accuracy [16, 17, 24]. Hence, to enable more flexible sparse pattern, we follow [19, 27] and define two kinds of heads as
@@
Following [19], for streaming heads, we directly use the local and sink tokens for sparse attention computation. For the retrieval heads, directly using all tokens for attention computation like [19] is inefficient. Therefore, we further develop a page-based dynamic selection method for retrieval heads to minimize the overhead of attention computation.
@@ IV-A1 Head Identification
As shown in Figure 5(a), we follow [19] and adopt an optimization-based method to identify two kinds of heads using the gating parameter $\alpha$. During training, we regard the retrieval head as full attention. The optimization can be formulated as
@@ IV-A3 Dynamic Sparsity of Retrieval Heads
retrieval heads dynamically select relevant tokens based on the current query. Inspired by [17, 24, 33], we group contiguous KV caches into pages for higher selection efficiency. We use the minimum and maximum boundary of each page to approximate the relevance score [17, 24]. Each page can be represented using metadata $\tau$ where
@@
To further minimize the overhead of dynamic selection, we share the page selection for near queries [27].
Memory consideration. Considering the memory limitation and hardware constraint, the increasing KV cache size during the decoding process is not friendly. Hence, we keep and update the accumulated attention scores for pages based on the computed relevance score [16] at each step, and then, evict one page with the lowest accumulated attention score when the KV cache size reaches the memory budget.
@@ IV-A4 Hardware Co-Design for Hybrid Sparse Attention
Streaming heads only attend to the sink and local tokens while retrieval heads are more complex: metadata needs first to be loaded into the logic die for $\min/\max$ computation, followed by the top-$k$ selection from the KV cache. Then, the selected KV cache is loaded into the logic die for attention computation. Following [22], retrieval heads also attend to sink and local tokens. As shown in Figure 6, we implement the following designs on hardware:
@@ IV-B2 Memory-Compute Co-Placement for Multi-Bank KV Cache
Load balancing introduces a new challenge, i.e., how to effectively distribute KV cache operations across multiple memory banks? Unlike the unified memory architecture of GPU, HB technology employs physically distributed memory where each bank maintains an independent storage resource. These challenges motivate us to simultaneously optimize both computation and memory allocation. Therefore, we propose the memory-compute co-placement that maps each token’s KV cache computation and their associated storage to the same bank. This architectural principle ensures localized processing of all KV cache operations for assigned tokens. When computing Softmax, we efficiently handle cross-bank communication following FlashAttention [34]. As long as the KV cache needs to be distributed across banks, we apply the memory-compute co-placement strategy.
@@ IV-C2 Mapping Heads of Different Sparsity
Our objective is to minimize the maximum distance between two kinds of heads since the workload share brings communication overhead. Figure 8(a) shows the difference between naive tiling without considering the communication distance and tiling with communication distance minimized. We model and solve this problem as a network maxflow problem [35] as illustrated in Figure 8(b).
@@ V-A1 Hardware Setups
We develop a cycle-level simulation framework to model the HB-based accelerator, integrating logic and memory dies vertically. The memory sub-system employs a multi-stack architecture consisting of 4 layers, with a macro bandwidth of 256 bits [11] and access energy 0.88pJ/bit [12]. Following [36], we build and model the General Matrix Multiplication (GEMM) accelerator consisting of 16 DCIM macros, each with 900 GOPS throughput and 24 TOPS/W energy efficiency. Detailed specifications are presented in Table II.
TABLE II: Hardware specification of H2EAL [11, 12, 36].
@@ V-A2 Software Setups
For LLM architectures, We evaluated the end-to-end performance on Mistral-7B, LLaMA2-7B, and LLaMA3-8B (quantized to 8-bit weights, 8-bit activations, and 8-bit KV)[37], which are popular models in edge scenarios. For evaluation datasets, the accuracy evaluation is conducted on LongBench [38] (up to 22k words) and Needle-in-a-Haystack (NIAH) [39] benchmarks.
@@ V-C Accuracy Evaluation
We evaluate the accuracy of our proposed H2EAL on the LLaMA3-8B [37] across the 21 datasets from LongBench [38] and the Needle-in-a-Haystack (NIAH) benchmark [39].
Figure 12: The accuracy comparison of H2EAL with full attention, StreamingLLM [22] and H2O [16] across multiple datasets in LongBench [38].

### v2

@@ I Introduction
Large language models (LLMs) have revolutionized various domains in field of generative artificial intelligence (AI) [Yi2024Survey]. While deployment on the cloud remains prevalent for mass user access, emerging trends show LLMs are deployed on intelligent edge platforms to address growing demands for localized functionality, personalized interactions, and data privacy [Durante2024AgentAI, Huang2022LanguageModels]. Edge-oriented LLM services prioritize small-batch processing while addressing growing demands for long-context applications such as long document analysis [Koh2022AnES], multi-turn dialogue [Zhang2025ASO], long chain-of-thought reasoning [Chen2025TowardsRE, Qiu2025PHYBenchHE]. Figure 1 demonstrates the trend in long-context processing and highlights hardware challenges in KV cache management.
Efficient LLM inference relies on the hardware to process two phases, i.e., prefill and decoding. During the prefill phase, LLM processes the input prompt, leveraging parallel processing to process multiple tokens. In contrast, during the decoding phase, LLM follows auto-regressive generation and only outputs one token at each step, which is more dependent on the memory due to the low data reuse. To meet both demands, near-memory processing (NMP)-based heterogeneous architectures have been proposed [Heo2024NeuPIMsNH, Yun2024DuplexAD, Park2024AttAccUT]. NMP combines standard computing units with memory-integrated accelerators to improve both compute-intensive and memory-heavy operations. However, traditional NMP approaches with in-die DRAM integration suffer from limited logic density and storage capacity [Devaux2019TheTP]. These challenges lead to lower throughput of attention computation and less effective key-value (KV) cache management as context lengths increase, ultimately affecting real-time performance in scenarios with long contexts and low batch sizes.
@@
The recent emerging hybrid bonding (HB) technology provides a promising alternative to conventional NMP architectures [Yue2024ExploitingSO, Niu2022184QPSW6L, Fujun2020ASE, Wuu20223DVT, h2llm]. It not only delivers substantial bandwidth with lower power consumption than HBM[Fujun2020ASE, Niu2022184QPSW6L], but also boosts NMP’s computational capacity by allowing processing engines to be implemented on the vertically stacked logic die. However, it exhibits distinctive features that differentiate it from other architectures. HB adopts a distributed-memory design, where the banks in memory dies are not directly interconnected, but connected to logic banks of high bandwidth with separate I/O and control units. This means that each bank operates largely independently, with a relatively limited storage capacity per bank.
Meanwhile, recent advancements in sparse attention methods [Zhang2023H2OHO, Tang2024QuestQS] have shown strong potential for edge inference in long-context scenarios. These approaches significantly reduce KV cache overhead with minimal impact on the model accuracy, becoming particularly beneficial for resource-constrained edge environments. However, previous methods are primarily designed for GPU-like systems, which rely on unified memory and enable seamless data sharing across all processing cores. These methods are not suitable for HB architectures due to the distributed memory. Additionally, different attention sparsity across heads has been observed in [Xiao2024DuoAttentionEL, Fu2024NotAH]. Therefore, allocating a unified sparse pattern for all heads is insufficient to fully leverage the sparsity.
@@ II-A HB Architecture for Edge-side LLM Inference
Deploying LLMs on edge devices for smart homes, autonomous driving, and IoT demands small-scale, real-time tasks to meet latency and power constraints [Durante2024AgentAI, Huang2022LanguageModels]. For long-context LLM inference where models handle extended lengths of sequence, edge devices face a unique challenge since the KV cache size becomes significant [Luohe2024KVCache]. Storing and accessing large KV cache require high memory bandwidth demands, which are difficult for traditional memory architectures, especially for edge devices with limited resources.
Emerging HB technology presents a promising solution by vertically stacking memory dies (e.g., DRAM) on top of logic dies and interconnecting them through high-bandwidth copper (Cu-Cu) direct bonding [Yue2024ExploitingSO, Niu2022184QPSW6L, Fujun2020ASE, Wuu20223DVT] as illustrated in Figure 3(a). This 3D integration enables substantial bandwidth and low power consumption, making HB a promising alternative to HBM and in-die NMP. However, despite its advantages, HB still faces challenges in managing memory and compute resources that are distributed spatially, especially for long-context inference where the KV cache becomes the dominant. Balancing the computational demands of LLMs with the memory bandwidth available in HB architecture is crucial for edge-side performance.
@@ II-B Sparse Attention Algorithm
Long-context LLM inference increases the computation and memory requirements induced by the KV cache, which becomes the prominent bottleneck. To alleviate this, sparse attention techniques have been proposed to reduce the number of attended tokens during each decoding step. These methods can be divided into static sparse attention (pattern-based) and dynamic sparse attention (input-based). Static sparsity such as the first few or local tokens uses pre-defined attention patterns to limit the number of tokens retained in the KV cache, which directly reduces memory usage but may risk losing important context [Xiao2023EfficientSL, Ge2023ModelTY]. In contrast, dynamic sparsity prunes less relevant tokens according to the input-based attention scores, offering better flexibility and utilization of the memory [Zhang2023H2OHO, Tang2024QuestQS, Chen2024ArkValeEG].
Hardware-based sparse attention has been explored to reduce memory usage and computational overhead in LLMs [Tu2023TranCIMFB, sadimm, Zhang2023H2OHO]. TranCIM [Tu2023TranCIMFB] adopts fixed-pattern sparsity based on the StreamingLLM algorithm [Xiao2023EfficientSL], restricting attention to local tokens and reducing data movement but yielding inflexible, suboptimal long-context performance. SADIMM [sadimm] employs top-k attention-based dynamic sparsity, improving adaptability but incurring hardware overhead for token ranking and risking accuracy loss[Tang2024QuestQS, Chen2024ArkValeEG]. H2O [Zhang2023H2OHO] evicts low-importance tokens yet underuses attention sparsity[Xiao2024DuoAttentionEL, Yang2025LServeEL], and remains bottlenecked by GPU memory bandwidth during decoding. As shown in Table I, these studies underscore the challenges in balancing sparse attention efficiency with hardware complexity, latency, and power consumption.
@@
Method TranCIM[Tu2023TranCIMFB] SADIMM[sadimm] H2O[Zhang2023H2OHO] H2EAL (This work)
@@ III-A Basic LLM Implementation on HB Architectures
Apart from regular multi-head attention (MHA), group-query attention (GQA) [Ainslie2023GQATG] and multi-query attention (MQA) [Shazeer2019FastTD] are proposed to alleviate the memory demand of MHA as illustrated in Figure 3(c). Different attention mechanisms are essentially the same as we focus on the KV head. Specifically, when centering around the KV head, once the KV cache is properly mapped and handled, the remaining weight modules can be allocated to the corresponding group of banks by tensor parallelism. However, implementing LLMs on the HB architecture on long contexts still faces the following challenges as shown in Figure 2.
@@ III-B Challenges
Challenge 1: Enormous KV cache overhead motivates the algorithm-hardware co-design. When handling long contexts, LLMs suffer from the increasing computation and memory requirements (including both memory capacity and bandwidth) associated with the KV cache, which becomes the prominent bottleneck for long-context LLM inference and exceeds the storage capacity of HB [Luohe2024KVCache] as illustrated in Figure 2. This challenge motivates us to leverage the inherent attention sparsity to compress the KV cache size. However, existing sparse attention methods either suffer from high accuracy degradation [Xiao2023EfficientSL] or introduce extra computation costs [Chen2024ArkValeEG, Tang2024QuestQS]. Meanwhile, recent studies observe that sparsity varies across different heads [Xiao2024DuoAttentionEL, Fu2024NotAH], providing us with larger optimization space for more flexible and extreme KV cache compression. These observations motivate us to co-design algorithm and hardware, including 1) an accurate head-wise sparse attention algorithm to achieve high sparsity without compromising accuracy; and 2) specialized hardware modules to better leverage the sparse attention with heterogeneous sparsity across heads.
Challenge 2: Varying sparsity in different heads leads to workload imbalance between banks. As described above, sparsity varies across different heads, shown in Figure 3(b). Instead of using uniform sparsity across heads, we apply different sparse patterns for different heads. However, such sparsity heterogeneity among heads leads to critical synchronization bottlenecks due to data dependencies - all heads must await completion of the most computationally intensive operation before progressing to subsequent layers. The challenge can be more severe due to the distributed memory architecture of HB architecture [Yue2024ExploitingSO, Niu2022184QPSW6L, Fujun2020ASE, Wuu20223DVT], which fundamentally differs from conventional unified memory systems. Hence, how to effectively balance workload distribution across memory-distributed banks for the hybrid sparsity pattern is a key question.
@@ IV-A Hybrid Sparse Attention with Algorithm-Hardware Co-Design
Transformer-based LLMs exhibit different sparse patterns in different heads [Xiao2024DuoAttentionEL, Yang2025LServeEL, Ge2023ModelTY, Fu2024NotAH]. Figure 4 shows the sparse patterns varies across different heads and layers. As shown in challenge 1 in Figure 2, some heads dynamically attend to the tokens based the the query while other heads mainly attend to the local and the first few tokens (termed as sink tokens) as observed in [Xiao2024DuoAttentionEL]. However, existing works set uniform sparse ratio for different heads, which is not sub-optimal for both efficiency and accuracy [Zhang2023H2OHO, Tang2024QuestQS, Chen2024ArkValeEG]. Hence, to enable more flexible sparse pattern, we follow [Xiao2024DuoAttentionEL, Yang2025LServeEL] and define two kinds of heads as
@@
Following [Xiao2024DuoAttentionEL], for streaming heads, we directly use the local and sink tokens for sparse attention computation. For the retrieval heads, directly using all tokens for attention computation like [Xiao2024DuoAttentionEL] is inefficient. Therefore, we further develop a page-based dynamic selection method for retrieval heads to minimize the overhead of attention computation.
@@ IV-A1 Head Identification
As shown in Figure 5(a), we follow [Xiao2024DuoAttentionEL] and adopt an optimization-based method to identify two kinds of heads using the gating parameter $\alpha$. During training, we regard the retrieval head as full attention. The optimization can be formulated as
@@ IV-A3 Dynamic Sparsity of Retrieval Heads
retrieval heads dynamically select relevant tokens based on the current query. Inspired by [Tang2024QuestQS, Chen2024ArkValeEG, zeng2025mpcache], we group contiguous KV caches into pages for higher selection efficiency. We use the minimum and maximum boundary of each page to approximate the relevance score [Tang2024QuestQS, Chen2024ArkValeEG]. Each page can be represented using metadata $\tau$ where
@@
To further minimize the overhead of dynamic selection, we share the page selection for near queries [Yang2025LServeEL].
Memory consideration. Considering the memory limitation and hardware constraint, the increasing KV cache size during the decoding process is not friendly. Hence, we keep and update the accumulated attention scores for pages based on the computed relevance score [Zhang2023H2OHO] at each step, and then, evict one page with the lowest accumulated attention score when the KV cache size reaches the memory budget.
@@ IV-A4 Hardware Co-Design for Hybrid Sparse Attention
Streaming heads only attend to the sink and local tokens while retrieval heads are more complex: metadata needs first to be loaded into the logic die for $\min/\max$ computation, followed by the top-$k$ selection from the KV cache. Then, the selected KV cache is loaded into the logic die for attention computation. Following [Xiao2023EfficientSL], retrieval heads also attend to sink and local tokens. As shown in Figure 6, we implement the following designs on hardware:
@@ IV-B2 Memory-Compute Co-Placement for Multi-Bank KV Cache
Load balancing introduces a new challenge, i.e., how to effectively distribute KV cache operations across multiple memory banks? Unlike the unified memory architecture of GPU, HB technology employs physically distributed memory where each bank maintains an independent storage resource. These challenges motivate us to simultaneously optimize both computation and memory allocation. Therefore, we propose the memory-compute co-placement that maps each token’s KV cache computation and their associated storage to the same bank. This architectural principle ensures localized processing of all KV cache operations for assigned tokens. When computing Softmax, we efficiently handle cross-bank communication following FlashAttention [Dao2022FlashAttentionFA]. As long as the KV cache needs to be distributed across banks, we apply the memory-compute co-placement strategy.
@@ IV-C2 Mapping Heads of Different Sparsity
Our objective is to minimize the maximum distance between two kinds of heads since the workload share brings communication overhead. Figure 8(a) shows the difference between naive tiling without considering the communication distance and tiling with communication distance minimized. We model and solve this problem as a network maxflow problem [Ford1956MaximalFT] as illustrated in Figure 8(b).
@@ V-A1 Hardware Setups
We develop a cycle-level simulation framework to model the HB-based accelerator, integrating logic and memory dies vertically. The memory sub-system employs a multi-stack architecture consisting of 4 layers, with a macro bandwidth of 256 bits [Yue2024ExploitingSO] and access energy 0.88pJ/bit [Niu2022184QPSW6L]. Following [Tsmc2021cim], we build and model the General Matrix Multiplication (GEMM) accelerator consisting of 16 DCIM macros, each with 900 GOPS throughput and 24 TOPS/W energy efficiency. Detailed specifications are presented in Table II.
TABLE II: Hardware specification of H2EAL [Yue2024ExploitingSO, Niu2022184QPSW6L, Tsmc2021cim].
@@ V-A2 Software Setups
For LLM architectures, We evaluated the end-to-end performance on Mistral-7B, LLaMA2-7B, and LLaMA3-8B (quantized to 8-bit weights, 8-bit activations, and 8-bit KV)[Lin2024QServeWQ], which are popular models in edge scenarios. For evaluation datasets, the accuracy evaluation is conducted on LongBench [longbench] (up to 22k words) and Needle-in-a-Haystack (NIAH) [kamradt2024llmtest] benchmarks.
@@ V-C Accuracy Evaluation
We evaluate the accuracy of our proposed H2EAL on the LLaMA3-8B [Lin2024QServeWQ] across the 21 datasets from LongBench [longbench] and the Needle-in-a-Haystack (NIAH) benchmark [kamradt2024llmtest].
Figure 12: The accuracy comparison of H2EAL with full attention, StreamingLLM [Xiao2023EfficientSL] and H2O [Zhang2023H2OHO] across multiple datasets in LongBench [longbench].